Suppose there are 100 firms in a perfectly competitive (price-taking) market and they each individually have the following average total costs:

Quantity 0 1 2 3 4 5 6
ATC - 200 140 110 105 104 110

         

If the market equilibrium quantity is 400 units, what is the market equilibrium price?

Group of answer choices

a $420

b $90

c $120

d $105
Answer #1

Answer : The answer is option b.

Here number of firms in the market is 100. For 400 units market quantities the each firm's quantity = 400 / 100 = 4 units.

TC (Total Cost) = ATC * Quantity

For firm's 3 units quantity level the TC = 110 * 3 = $330.

For firm's 4 units quantity level the TC = 105 * 4 = $420.

MC (Marginal cost) for firm's 4th unit = TC of 4 units - TC of 3 units = 420 - 330 = $90.

For perfectly competitive firm the market price is equal to MC. So, here the market Price = MC = $90. This means that the market equilibrium price is $90.

Hence except option b other options are not correct. Therefore, option b is the correct answer.